OpenStudy (anonymous):

The sum of the interior angles of a 4 sided, regular polygon is 720 degrees. Select one: True False

OpenStudy (acxbox22):

false

OpenStudy (acxbox22):

sum of interior angles= 180(n-2) where n is the number of sides 180(4-2)=180(2)=360 degrees